OpenStudy (anonymous):

Using dimensional analysis, convert 981 dL to kL. (Hint: 10 dL = 1 L; 1,000 L = 1 kL) A. 9.81 × 10^6 kL B. 9.81 × 10^3 kL C. 9.81 × 10^1 kL D. 9.81 × 10^–1 kL E. 9.81 × 10^–2 kL

OpenStudy (anonymous):

Can anybody please help?

OpenStudy (anonymous):

use the same way to solve this question as you did the previous one... 1kL = 1000L = 1000 * (10 dL) as 1 L = 10 dL this means 1kL = 10000 dL i think now you can find out your answer,,

OpenStudy (anonymous):

is 9.81 / 1000 the correct step?

OpenStudy (anonymous):

why you wrote 9.81?

OpenStudy (anonymous):

sorry 981 * not 9.81

OpenStudy (anonymous):

gd.. and are you not missing a zero in the denominator??? i mean you wrote 981/1000 i.e. 3 zeroes but there should be 4 zeroes.. find out why...

OpenStudy (anonymous):

if 10000 dL = 1kL then, 1dL = 1/10000 kL okay with this? and 981 dL = 981/10000 kL keep it simple.. work it out you willl get the answers to all the questions of this sort

OpenStudy (anonymous):

Thanks a lot!!! Is E the correct answer?

OpenStudy (anonymous):

@ANISH95

OpenStudy (anonymous):

yes absolutely right ..! its E...

OpenStudy (anonymous):

Thank you so much for everything!

OpenStudy (anonymous):

:-) no probemm